# On the Feasibility of Hijacking MLLMs’ Decision Chain via One Perturbation

### v1

@@ 4 Latent Space Analysis
Empirical Observations in the Latent Space. We conduct experiments by searching for an SAUP on Llava-1.5-7B [22] using images from five distinct classes of ImageNet [9]. For each class, we randomly select 20 images and assign a unique word as the target label. We then extract and visualize the latent features from the model’s penultimate layer.
As shown in Figure 4, our investigation of the latent space yields three progressive observations:
Separation of Clean and Perturbed Features. First, we observe a clear separation between the features of clean and perturbed images. To visualize this, we perform principal component analysis (PCA) on the extracted features. As shown in Figure LABEL:fig:pca, the features of perturbed images are clustered together, far from the widely dispersed features of the original clean images. This indicates that the SAUP imposes a dominant influence on the features.
Semantic Separation within Perturbed Features. Second, while the perturbed features are grouped together, they are not monolithic. A closer look reveals that they form distinct sub-clusters based on their semantics. Figure LABEL:fig:pca3d visualizes the perturbed features in a 3D view, showing that features corresponding to different semantics occupy separate spaces. This suggests the perturbed features are deflected by their original semantics.
Alignment of Semantic Clusters with Target Labels. Finally, we observe that these semantically segregated clusters are precisely aligned with their designated targets. We compute the average output confidence for each target, as illustrated in Figure LABEL:fig:alignment. The high values along the diagonal confirm that perturbed images from a specific semantic class consistently elicit a high confidence score for their predefined target. This demonstrates that, after semantic deflection, the perturbed features are well aligned with their predefined targets in the latent space.
Figure 5: Illustration of the underlying mechanism of the Semantic-Aware phenomenon. Given three image sets $\mathcal{V}_{0},\mathcal{V}_{1},\mathcal{V}_{2}$ with distinct semantic contents, and an all-zero pixel image serving as an anchor. The SAUP maps all image features to a distant region in the latent space, while the respective semantic directions (SDs) of $\mathcal{V}_{0},\mathcal{V}_{1},\mathcal{V}_{2}$ cause slight deflections in these features, guiding them towards alignment with predefined targets.
Explanation for the Semantic-Aware Phenomenon. Based on these observations, we propose an explanation for the semantic-aware nature of SAUPs. We posit that: Image features under the SAUP are dominated by the perturbation and slightly deflected according to the image semantics.
This mechanism can be understood with a simple model, as illustrated in Figure 5. Let’s consider image sets $\mathcal{X}_{0},\mathcal{X}_{1},\mathcal{X}_{2}$ with distinct semantics. The SAUP introduces a common directional vector to all images, pulling their features into a specific region of the latent space, far from their “clean” locations. We can conceptualize this “pure” perturbation feature by applying the SAUP to an anchor image with no semantic content (e.g., an all-zero image).
Meanwhile, the inherent semantic content provides a smaller, secondary vector that deflects the feature away from this pure perturbation point. Since different classes (e.g., $\mathcal{X}_{0}$ vs. $\mathcal{X}_{1}$) have different semantic directions, the perturbed features are deflected into separate spatial clusters.
Therefore, the optimization algorithm can be viewed as searching for a point in the latent space such that the spatial clusters formed by the deflected features are aligned with the predefined targets. By aligning different perturbed feature clusters with their corresponding targets, the perturbation appears to acquire the ability to “actively” recognize the semantics of images and assign targets accordingly.
Figure 6: Attack success rates versus the number of targets (#targets) on the ImageNet. The frame refers to adding a narrow border around an image, while the corner refers to attaching patches to the four corners of an image.
5 Experiments
@@ 5.1 Experimental Setup
Datasets. We evaluate the effectiveness and generalization of SAUPs on two datasets.
(1) ImageNet. To assess the effectiveness of SAUPs on coarse-grained semantics, we employ the ImageNet [9], which serves as a widely used benchmark in prior adversarial studies [10, 7]. We first randomly select $C$ classes from ImageNet. For each class, 50 images are used as the training set (i.e., $50\times C$ images in total), and another 20 non-overlapping images are used as the test set. The target for each class is randomly selected from the remaining categories and is represented by a single word; sentence-level targets are evaluated in Section 5.3.
(2) RIST. To assess the effectiveness of SAUPs on fine-grained semantics, we annotate and evaluate the RIST dataset. RIST consists of similar frames extracted from real-world videos, covering two scenarios: AutoDriving and RoboTasking with 5 and 2 classes, respectively. Each class contains 10 images, as well as a manually annotated word serving as the target. The 5 images are used for training, and the remaining 5 images are used for testing.
For collecting RIST, our annotators meticulously identify segments that contain potential risks and exclude repetitive or semantically meaningless segments. Each category within the RIST dataset contains sequences centered around a specific event, such as robots interacting with kitchen utensils or cars passing through intersections with pedestrians and warning signs.
To ensure each sequence captures a unique event while maintaining diversity between different sequences, our annotators review more than 1 TB of video frames and manually filter out highly similar images. We ensure that every image in the RIST dataset is independently cross-validated by two annotators to guarantee quality and consistency. More details can be found in the Appendix.
@@
Models. Our experiments involve three representative MLLMs. (1) LLaVA-1.5-7B (Llava) [22] is a widely used open-source MLLM that has been extensively used as a benchmark in prior studies [23, 17]. (2) Qwen2.5-VL-7B-Instruct (Qwen) [3] is a well-known model that incorporates advanced instruction-following abilities. (3) InternVL3-8B (Intern) [42] is designed for strong general-purpose vision and language understanding, with competitive performance across multiple vision-language tasks [11].
@@
Following [23], we default to fixing the user input prompt $\mathbf{p}$ as Describe this image, and utilize greedy decoding for both MLLMs. We also evaluate attack performance under varying input prompts in the Appendix.
@@
Perturbation Constraints. Consistent with prior work [23], we focus on two types of constrained adversarial perturbations: frame and corner, where the frame width is set to $6$ pixels, and each corner patch has a size of $20\times 20$ pixels unless otherwise specified  [^2] These perturbation constraints are defined for an image of size $300\times 300$ [31, 5]. For fair comparison across models, we linearly scale constraints according to the specific input size required by each model. . We also involve more constraint settings in the Appendix.
Evaluation Metric. Following [23, 31], we report the Attack Success Rate (ASR) as the evaluation metric. ASR is a simple yet stringent measure, as it requires the model’s predictions to exactly match the predefined target sequences.
Given a dataset $\mathcal{D}=\left\{\left(\mathbf{x^{\prime}}_{i},\mathbf{t}_{i}\right)\right\}_{i=1}^{|\mathcal{D}|}$, where each $\mathbf{x^{\prime}}_{i}$ is a perturbed image and $\mathbf{t}_{i}$ is its corresponding target (a word or a sentence), the ASR is formally defined as:
$\text{ASR}=\frac{1}{|\mathcal{D}|}\sum_{i=1}^{|\mathcal{D}|}\mathbb{I}\left[f(\mathbf{x}_{i}^{\prime},\mathbf{p})=\mathbf{t_{i}}\right]$ (10)
where $\mathbb{I}[\cdot]$ is the indicator function that returns 1 if the condition is true, and 0 otherwise.
Considering the model may require outputting many tokens to convey meaningful content, these tokens must match the target sequence in both content and order. This strict metric captures the ability of SAUPs to exert precise control, which is essential for downstream applications.
Implementation Details. The experiments are conducted using seven NVIDIA RTX 3090 GPUs. Consistent with previous work [23], each SAUP is generated over 500 epochs. We employ the Momentum optimizer [10] with a momentum of 0.9, a learning rate of 0.01, and set the margin loss weighting factor to 1. During inference, greedy decoding is adopted, that is, the token with the highest confidence is selected at each step.
@@
5.2 Vulnerability Analysis of MLLMs
5.2.1 Results on the ImageNet
We assess the effectiveness of SAUPs with different numbers of targets (#targets) on ImageNet. As shown in Figure 6, we have the following findings:
(1) MLLMs exhibit significant vulnerability to SAUPs. When misleading five targets, we achieve the ASR over 63% and 45% with frame and corner constraints, respectively. Notably, when attacking 9 targets under the frame constraint, we still achieve ASRs of 42% on Llava, 31% on Qwen, and 48% on Intern. These results demonstrate that SAUPs can effectively aware the semantics of input images and steer the model toward distinct targets.
(2) The ASR exhibits a decreasing trend as #targets increases. We evaluated SAUP’s performance with one to nine targets and observed a general decline in ASR as the number of targets grew. Among the different model types, Intern demonstrated the greatest susceptibility, as its ASR curve decreased more gradually than the others.
(3) The perturbation constraint plays a critical role in determining ASR. For Llava and Qwen, the ASR of corner-based perturbations drops to 0 when the number of targets reaches eight. This phenomenon can be attributed to the disparity in the number of modifiable pixels, 1,600 for corners versus 7,056 for the frame. In contrast, corner-based perturbations achieve comparable ASR with a small number of targets because their perturbations are spatially concentrated, allowing them to effectively manipulate features.
5.2.2 Results on the RIST
In this subsection, we evaluate the performance of SAUPs on the RIST dataset. Unlike traditional classification datasets, RIST focuses on fine-grained semantic understanding and covers two scenarios: RoboTasking (2 categories) and AutoDriving (5 categories). We summarize the ASR for each trajectory, as shown in Figure 8.
The results demonstrate that we can generate effective SAUPs based on fine-grained semantic image clustering. On the test sets of the RoboTasking and AutoDriving scenarios, we achieve an average ASR of 72% and 62%, respectively. These results highlight the flexibility and efficiency of SAUP generation, revealing that attackers can craft highly effective perturbations from only five images, thereby posing a significant security risk.
The overfitting phenomenon occurs during the SAUP generation. For example, we achieve a 100% ASR on the training set for Intern for both scenarios, yet the average ASR drops to 73% and 61% on the test set, respectively. This overfitting arises from the limited training set of RIST, which involves at most 50 images for training, making it easy for the SAUP to overfit. We think this issue can be mitigated by increasing the size of the training set and incorporating multiple image augmentation strategies.
@@ 5.2.2 Results on the RIST
Method Baseline w/o NSO w/o MarginLoss Default Baseline w/o NSO w/o MarginLoss Default
@@
5.3 Evaluation of Sentence Control
Precise control over sentence-level outputs expands the application of image hijacks, enabling scenarios such as spreading misinformation and executing malicious commands. This subsection evaluates the SAUPs’ ability to control sentence-level output by varying the number of words per target (#words per target). In other words, the SAUP is expected to compel the MLLM to generate sentences that exactly match the predefined targets for different images.
We evaluate the attack performance under the frame constraint by using up to 15 words for 2 targets and up to 4 words for 5 targets. As illustrated in Figure 7, we can have the following observations:
SAUPs are effective in controlling MLLMs to generate sentences. Impressively, when each target consists of 10 words under 2 targets, we still achieve an average ASR of 83% on the test set. When the number of targets increases to 5, SAUPs are still able to control 4 words per target, achieving 54% average ASR on the test set.
The Qwen and Intern are highly susceptible to control by SAUPs, enabling the precise generation of sentences. For example, when reaching 15 words under 2 targets, we achieve 78% ASR on Qwen and 83% ASR on Intern. This suggests that these models are heavily impacted by incorrect tokens in the preceding generation, which in turn makes the subsequent sequence more prone to being erroneous.
Figure 9: Loss curve during SAUPs optimization for the Qwen with 5 targets. The absence of NSO causes the loss to remain consistently high and prevents effective reduction during optimization. Removing the margin loss leads the optimization process to converge to local optima.
@@ 5.4 Ablation Studies
This subsection presents ablation studies to analyze the individual contributions of the two components in the proposed method, i.e., the normalized space optimization and the semantic separation optimization.
We first introduce two variants of the SORT: (1) w/o NSO, which removes the normalized space optimization module and directly optimizes SAUPs in the $[0,1]$ pixel space; (2) w/o MarginLoss, which eliminates the margin loss function and relies solely on the cross-entropy loss for optimization. We also adapt the existing image hijack [4] for multiple targeted scenarios to serve as a baseline.
We conduct experiments with #targets set to 2 and 5 on these variants. As shown in Table 1, the default method achieves the highest ASR across most settings, indicating the effectiveness of both components. We also visualize the global loss and margin loss on Qwen with 5 targets in Figure 9. In the following, we provide an in-depth analysis of two components, respectively.
The NSO module provides the foundation for optimization, ensuring the effectiveness of the SORT. Specifically, Table 1 shows that the absence of NSO leads to an average 49.5% decrease in ASR on the test set. Figure 9 further indicates that, without NSO, the loss becomes trapped in a plateau, making it difficult for the optimization to converge to an optimal solution. These results suggest that NSO is a fundamental component in the optimization of SAUPs.
The semantic separation module improves the effectiveness of the SORT, particularly as #targets increases. As shown in Table 1, the w/o MarginLoss variant achieves a comparable ASR to the default setting when #targets is two, as the optimization difficulty remains low and cross-entropy loss alone is sufficient to guide the attack. However, as #targets increases, the margin loss yields an average ASR gain of 6%. Figure 9 also shows that although the variant without margin loss converges faster initially, it ultimately tends to get trapped in a local optimum.
@@ 6 Conclusion
This paper demonstrates the feasibility of hijacking the decision chain of neural networks through a single perturbation, which compels the model to precisely output different predefined targets based on the input semantics. To address the optimization challenges posed by the alignment of input semantics and predefined targets, we introduce SORT, an effective algorithm that incorporates a normalized space optimization and a semantic separation strategy. We further introduce RIST, a newly annotated dataset to evaluate the attack performance under fine-grained semantic granularity.
We perform an in-depth analysis of perturbed and clean image features in the latent space to explain the underlying reasons for the semantic-aware ability. We conduct extensive experiments on both ImageNet and RIST datasets on three representative MLLMs and two perturbation constraints. The results demonstrate the effectiveness and generalization capability of SAUPs. For future work, we intend to develop robust and transferable SAUPs that can be deployed in the physical world. We hope this paper can inspire the discovery of novel security threats in neural networks.

### v2

@@ 5 Validation of Geometric Mechanics
We conduct experiments using Llava-1.5-7B (Liu et al., 2023), selecting images from five distinct classes of ImageNet (Deng et al., 2009) and assigning a unique target label $t^{(c)}$ to each class. We visualize the penultimate layer features to corroborate our geometric propositions.
5.1 Validation of the Dominant Shift
Observation: As illustrated in LABEL:fig:pca, we observe a significant separation between the clusters of clean images and perturbed images in the latent space. To confirm that this shift is intrinsic to the perturbation itself, we feed a “pure perturbation” (an all-zero image added with $\delta$) into the model. As shown in LABEL:fig:pca, this point lies at the centroid of the perturbed image clusters.
Conclusion: This observation validates the existence of the Dominant Shift $\mu_{\delta}=\phi(\delta)$, derived in Equation 3.
5.2 Validation of Semantic Deflection
Observation: While perturbed features are clustered away from clean data, LABEL:fig:pca3d reveals that they are not monolithic. Instead, they form distinct, separable sub-clusters corresponding to their original semantic categories. This formation of these distinct sub-clusters proves that the Semantic Deflection $\nu_{c}$ can identify the semantic differences of inputs, and deflect to different directions.
Conclusion: This confirms the role of the Semantic Deflection $\nu_{c}=J_{\delta}\cdot x^{(c)}$, derived in Equation 3.
5.3 Validation of Alignment with Targets
Observation: LABEL:fig:alignment presents the confidence heatmap, which shows that the perturbation results in a significantly higher confidence along the diagonal. This indicates that the direction of semantic deflection is not stochastic. Instead, it forces the embeddings of input images to align precisely with their respective proxy target embeddings, thereby triggering the predefined targets.
Conclusion: This indicates that deflected embeddings are aligned with corresponding proxy target embeddings, derived in Equation 6
Figure 5: SAUP cases generated on a test image trajectory within RIST. We optimize and apply a single SAUP to these test images. Although different categories share similar semantic content (such as motorway and sidewalk), the perturbation successfully acts as a “semantic router,” forcing MLLM to output different targets based on the content of the current frame.
6 Performance Evaluation
@@ 6.1 Experimental Setup
Datasets.
We evaluate the effectiveness and feasibility of SAUPs on two datasets. Each dataset contains a train set for optimizing perturbations and a non-overlapping test set for evaluating the generalization ability of the perturbations.
(1) RIST (Fine-grained Semantics). Recall from Remark 3.5 that fine-grained semantics are more difficult to attack. To evaluate the performance of semantic-aware hijacking under this task, we annotate the Real Image Sequence Trajectories (RIST) dataset.
Unlike standard classification benchmarks, RIST consists of sequential frames extracted from continuous videos (AutoDriving and RoboTasking scenarios), serving as a stringent stress test. More details are depicted in Appendix A.
We leverage the Gemini-2.5-pro (DeepMind, 2025) to randomly assign a specific, contextually relevant action (e.g., “accelerate” at a red light) as the target $t^{(c)}$ for each semantic cluster $c$, based on the scenario’s safety constraints.
(2) ImageNet (Coarse-grained semantics). We employ the ImageNet (Deng et al., 2009) to assess the effectiveness on coarse-grained semantics. The target for each category is randomly selected from the remaining classes and completed into a full sentence using the Gemini.
@@
Models. Our experiments involve three representative MLLMs: (1) Llava-1.5-7B (Llava) (Liu et al., 2023), a standard open-source benchmark; (2) Qwen2.5-VL-7B-Instruct (Qwen) (Bai et al., 2025), known for advanced instruction-following; and (3) InternVL3-8B (Intern) (Zhu et al., 2025), a strong general-purpose model.
@@
Consistent with prior work (Lu et al., 2024), we default to fixing the user input prompt $\mathbf{p}$ as Describe this image, and utilize greedy decoding for both MLLMs. We evaluate varying prompts in Appendix B.
@@
Perturbation Constraints. We focus on two types of constrained adversarial perturbations: frame and corner, where the frame width is set to $6$ pixels, and each corner patch has a size of $20\times 20$ pixels unless otherwise specified. We also involve more constraint settings in Appendix C.
Evaluation Metric. We report the Attack Success Rate (ASR), which requires the output tokens to match the target sequence in both content and order.
Figure 7: SAUP cases of precise control over long-text generation. On the test set, we demonstrate that a single SAUP can bind complex, long textual targets to specific visual concepts. For instance, “truck” triggers a specific long sentence, while “mountain” triggers a crayfish description, verifying the model’s ability to maintain precise control over extended generation sequences.
@@
6.2 Feasibility of Semantic-Aware Hijacking
This section aims to verify the feasibility of semantic-aware hijacking under different semantic granularities.
6.2.1 Hijacking Fine-Grained Semantics
We first demonstrate the “Semantic-Aware” capability of SAUPs through a visualization on the RIST dataset. Figure 5 shows images from the test set of RIST, which are extracted from a continuous video. The adversary optimizes and applies a single universal perturbation to this continuous image trajectory. Although different categories share similar semantic content (such as motorway and sidewalk), the SAUP successfully amplifies these subtle semantic variations in the latent space. As MLLM processes image trajectory, the perturbation acts as a semantic router, forcing the model to output distinct, attacker-defined, and dangerous commands such as “accelerate”, “merge”, or “steer” at precise moments. This confirms that SAUPs can effectively hijack multiple stateless decisions by binding specific visual semantics to predefined targets.
To systematically evaluate this threat, we evaluate the ASR across the entire RIST dataset, covering two scenarios: RoboTasking (2 targets) and AutoDriving (5 targets). As shown in Figure 6, SAUPs achieve an average ASR of 72% on the RoboTasking test set and 62% on the AutoDriving test set. It is worth noting that we observed an overfitting phenomenon. For instance, on the Intern model, the ASR reaches 100% on the training set but drops to 61% on the test set for the AutoDriving scenario. This performance gap is attributed to the limited size of the RIST training set (50 images). Nevertheless, these results demonstrate that even with limited data, a single perturbation can learn to distinguish fine-grained semantics and successfully commandeer the model’s decision process across multiple targets.
Figure 9: Attack success rates versus the number of targets (#Targets).
Figure 10: Loss curve during optimization for Qwen.
@@ 6.2.1 Hijacking Fine-Grained Semantics
Method Baseline w/o NSO w/o SSO Default Baseline w/o NSO w/o SSO Default
@@
6.2.2 Precise Control over Long Output
Beyond simple semantic routing, an adversary may need to inject complex payloads, such as specific sentences or malicious scripts. Figure 7 visualizes the capability of SAUPs to generate long, coherent text sequences on the ImageNet dataset. We assign semantically unrelated target sentences to different image classes. For example, the “truck” is bound to a poetic description, while “mountain” triggers a factual statement about crayfish. The results show that the perturbation effectively and precisely controls the generation of long sequences by the MLLM.
We comprehensively evaluate the effectiveness of this control by varying the sequence length (Number of Words per Target). As illustrated in Figure 8, SAUPs exhibit strong capacity in controlling long outputs. When each target consists of 10 words under 2 targets, we still achieve an average ASR of 83% on the test set. When the number of targets increases to 5, SAUPs are still able to control 4 words per target, achieving 54% average ASR on the test set. The Qwen and Intern are highly susceptible to control by SAUPs, enabling the precise generation of sentences. For example, when reaching 15 words under 2 targets, we achieve 78% ASR on Qwen and 83% ASR on Intern. This suggests that these models are heavily impacted by incorrect tokens in the preceding generation, which in turn makes the subsequent sequence more prone to being erroneous.
6.3 Information Capacity of Single Perturbation
While Theorem 3.4 establishes the theoretical feasibility bound, this condition primarily governs the feasibility of the “worst-case” pair mapping. However, as the number of targets increases, the challenge shifts from the magnitude of the expansion to the capacity of the perturbation to encode multiple distinct mappings simultaneously. In this subsection, we evaluate the capacity limit of a single perturbation by increasing the number of targets  [^3] Each target is represented by a single word, using 50 images for training and 20 non-overlapping for testing. on ImageNet.
As illustrated in Figure 10, we observe a general decline in ASR as the number of targets increases. For example, on Qwen, the ASR under the frame constraint drops from 93% (2 targets) to 31% (9 targets). We attribute this phenomenon to the limited information capacity of the perturbation. Recall Equation 5 that $\|J_{\delta}\|_{2}$ can quantify the theoretical mapping capability for the perturbation $\delta$. Since $\delta$ is spatially constrained (i.e., restricted to a frame or corners), its limited number of trainable pixels creates an information bottleneck. As the number of targets grows, it is increasingly difficult for a single $\delta$ to encode all required decision boundaries within the finite pixel budget.
This “information capacity” hypothesis is supported by the performance disparity between Frame and Corner constraints. The Frame perturbation occupies 7,056 pixels, providing a larger parameter space to shape the Jacobian $J_{\delta}$. Consequently, it maintains a relatively high ASR (e.g., 42% on Llava for 9 targets). In contrast, the Corner perturbation contains only 1,600 pixels. Due to this severe capacity restriction, the ASR for Corner perturbations collapses rapidly, dropping to 0% on Llava and Qwen when the number of targets reaches 8. Furthermore, the ASR can be significantly improved by slightly increasing the perturbed area (see Appendix C), further validating this hypothesis.
@@ 6.4 Ablation Studies
In this subsection, we disentangle the contributions of the two core components of SORT: Normalized Space Optimization (NSO) and Semantic Separation Optimization (SSO). We compare the default SORT with two variants: (1) w/o NSO, which optimizes directly in the $[0,1]$ pixel space; and (2) w/o SSO, which removes the margin loss $\mathcal{L}_{Margin}$ and relies solely on Cross Entropy. In addition, we applied an image hijacking method (Bailey et al., 2023) to this multi-target scenario as a baseline for comparison.
NSO: The Bedrock of Optimization. Although NSO introduces a simple change of variables (mapping from pixel space to a normalized unconstrained space), it serves as the fundamental prerequisite for our attack’s success.
Direct optimization in the pixel space is notoriously unstable for generating universal perturbations, often leading to loss plateauing. As evidenced in the left part of Figure 10, the loss curve of the “w/o NSO” variant remains consistently high and fails to descend, indicating that the optimizer is trapped in a poor local minimum.
This optimization failure results in a catastrophic performance drop. As shown in Table 1, removing NSO causes the ASR to plummet by an average of 49.5% on the test set. Specifically, on Qwen with 5 targets and Frame constraint, the ASR drops from 66% to near zero without NSO. These results confirm that NSO is not merely a trick, but a necessary condition to make the geometric alignment of $J_{\delta}$ mathematically solvable.
SSO: The Booster for Multi-Targeting. The SSO improves the effectiveness of the SORT, especially in complex scenarios. As shown in Table 1, the w/o SSO variant achieves a comparable ASR to the default setting when the number of targets is two, as the standard Cross-Entropy loss is sufficient to separate a few classes. However, as the semantic crowding increases, the necessity of SSO becomes apparent, yielding an average ASR gain of 6%. Figure 10 further reveals that while the optimization without SSO converges quickly in the early stages, it gets trapped in a local optimum. In contrast, the default method continues to optimize the margin, pushing the perturbed features toward the centers of their respective target clusters, and eventually achieves a smaller overall loss and better generalization.
@@ 7 Conclusion
This paper investigates a novel and severe threat to semantic-aware hijacking in MLLMs. We demonstrate that a single SAUP can effectively act as a “semantic router,” compelling the model to precisely output distinct, attacker-defined targets conditioned solely on the input semantics.
To achieve this, we introduce an optimization algorithm (SORT), and a dataset featuring fine-grained semantic annotations (RIST), which serves as a simplified realistic threat model for autonomous driving and robotic tasks.
We provide a comprehensive understanding of the underlying mechanism of SAUPs. We theoretically analyze the geometric feasibility bound and empirically validate the “Dominant Shift” and “Semantic Deflection” phenomena in the latent space. Extensive experiments across three representative MLLMs confirm the feasibility of SAUPs under various settings, including different semantic granularities, control lengths, and target diversities.
For future work, we intend to bridge the gap between digital simulation and reality by developing robust SAUPs resilient to physical conditions (e.g., brightness changes).
Impact Statement
This paper exposes Semantic-Aware Hijacking, a novel MLLM vulnerability where a single perturbation can semantically route multiple inputs to different malicious targets. The significance of this vulnerability lies in its capability to trigger cascading consequences by hijacking a chain of decisions. Given the rapid emergence of MLLM-based agents, such as those designed for automated computer and smartphone interaction, the potential for exploitation is vast. For instance, a malicious frame on a webpage that coerces an agent first uploads a user’s private data and then executes a “submit” command to finalize the breach.